# AI-driven Java Performance Testing: Balancing Result Quality with Testing Time

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXcopyright: noneconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software performanceccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Finding this balance becomes particularly difficult in the Java context, where software execution undergoes an initial warm-up phase due to just-in-time compilation (Barrett et al., 2017). During this phase, the Java Virtual Machine (JVM) performs a wide range of optimizations, leading to fluctuations in performance behavior (as shown in Fig. 1) that may adversely affect result quality (Georges et al., 2007; Traini et al., 2022). To mitigate this issue, it is common practice to conduct a number of “warm-up iterations”, with the sole goal of warming up the JVM, before starting to collect performance measurements.
@@
  • We make publicly available the source code of our framework and the pretrained TSC models to aid future studies [^1] Our replication package, including the dataset, source code and pre-trained TSC models, is publicly available at https://github.com/SpencerLabAQ/AIPT..
@@ 2.1. Java Microbenchmarking
Microbenchmarking is a type of unit-level performance testing, commonly used in Java context (Leitner and Bezemer, 2017). A microbenchmark repeatedly executes a small portion of code, such as a Java method, while gathering measurements of its execution time. Due to just-in-time compilation, microbenchmarks are subject to performance variability in the first phase of their execution, also known as warm-up. During this phase, the JVM detects frequently executed loops or methods, and it dynamically compiles them into optimized machine code. After the completion of the warm-up phase, the microbenchmark is said to be executing at a steady-state of performance(Barrett et al., 2017).
@@
State-of-practice (SOP)
Practitioners usually predefine a fixed number of warm-up iterations based on their domain expertise. To configure warm-up iterations, they typically rely on Java Microbenchmark Harness (JMH) [^2] https://github.com/openjdk/jmh, the de-facto standard for building, configuring, and running Java microbenchmarks. JMH allows to configure the number of warm-up iterations directly in the microbenchmark’s source code, using Java annotations, as shown in Listing 1. Nevertheless, prior work has shown that using a fixed number of iterations may often produce suboptimal estimates of the warm-up phase (Georges et al., 2007; Laaber et al., 2020; Traini et al., 2022).
State-of-the-art (SOTA)
Researchers have developed techniques that can dynamically stop warm-up iterations at run-time. Georges et al. (Georges et al., 2007) introduced a first such approach, using a preset threshold on the coefficient of variation (Everitt and Skrondal, 2010) to determine the end of the warm-up phase. However, subsequent studies have identified Georges et al.’s heuristic as both inaccurate (Kalibera and Jones, 2013) and unrealistic (Laaber et al., 2020; Traini et al., 2022) in practical scenarios. Another technique was recently proposed by Laaber et al. (Laaber et al., 2020). Similarly to Georges et al.’s heuristic, it leverages statistical metrics to estimate the end of the warm-up phase. At each microbenchmark iteration, this technique checks whether adding more performance measurements are likely to change the distribution of performance measurements, and it uses this information to dynamically stop warm-up iterations at runtime. Laaber et al. provide three different variants of this technique, based on distinct statistical metrics, namely coefficient of variation (Everitt and Skrondal, 2010) (COV)  [^3] Both the approaches of Georges et al. (Georges et al., 2007) and Laaber et al. (Laaber et al., 2020) use the CV, but they apply it in different ways. Georges et al. use a fixed threshold on CV, whereas Laaber et al. verify whether the addition of more measurements changes the CV within a specified threshold., relative confidence interval width (Davison and Hinkley, 1997; Kalibera and Jones, 2012) (RCIW), and Kullback-Leibler divergence (Kullback and Leibler, 1951; He et al., 2019) (KLD). Prior work has shown that Laaber et al.’s technique is more accurate than the SOP in estimating the end of the warm-up phase (Traini et al., 2022).
@@ 3.4. Application
Our framework employs TSC models to dynamically predict when the microbenchmark reaches a steady-state of performance. As the microbenchmark execution progresses, the framework continuously evaluates the incoming performance measurements using the TSC model. Once the model detects the achievement of the steady-state, the framework promptly halts the microbenchmark execution and returns the current set of measurements for performance evaluation. Fig. 3 presents a snapshot of this procedure using an illustrative scenario. In this scenario, the microbenchmark has completed 3 warm-up iterations, with the current measurement window ranging from the 4th to the 7th iteration. Before proceeding, our framework queries the TSC model, forwarding the current set of measurements to determine if they correspond to a steady-state of execution. If the model predicts that the measurements are stable, the microbenchmark execution is halted, and these measurements are returned for performance evaluation. Conversely, if the model classifies them as unstable, then the framework makes the test running another iteration, and it shifts the measurement window by one iteration. This process repeats at each iteration until the model identifies a stable segment (or until a predefined maximum number of iterations is achieved).
@@ 5.1. RQ1: To what extent can TSC models accurately classify stable and unstable measurements?
Approach. We extract the number of warm-up iterations specified by the developers for each microbenchmark using a method similar to (Traini et al., 2022). We then compare the WEE of our framework with that of the SOP for each time series that reaches a steady-state. For this comparison, we employ the Wilcoxon signed-rank test (Wilcoxon, 1945) along with two measures of effect size: the Vargha-Delaney $\hat{A}_{12}$ (Vargha and Delaney, 2000) and the matched pairs rank biserial correlation $r$ (Kerby, 2014). In our context, the $\hat{A}_{12}$ measures the proportion of pairs where the WEE of our framework is lower than that of the developers. An $\hat{A}_{12}$ value $>0.5$ indicates that our framework provides a more accurate estimation of the warm-up phase than the SOP. The matched pairs rank biserial correlation $r$ represents the difference between the proportion of favorable and unfavorable evidence; in our case, favorable evidence indicates a lower WEE for the framework. Thus, an $r>0$ indicates that our framework performs better than the SOP.
@@ 8. Conclusion
In this paper, we propose and study an AI-based framework to dynamically stop warm-up iterations in Java performance tests. The results show that, when integrated with certain TSC models, our framework delivers result quality comparable to the state-of-practice while drastically reducing the testing time. Furthermore, we find that the framework improves both the result quality and testing time in most of the investigated state-of-the-art techniques.

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: 39th IEEE/ACM International Conference on Automated Software Engineering ; October 27-November 1, 2024; Sacramento, CA, USAbooktitle: 39th IEEE/ACM International Conference on Automated Software Engineering (ASE ’24), October 27-November 1, 2024, Sacramento, CA, USAdoi: 10.1145/3691620.3695017isbn: 979-8-4007-1248-7/24/10copyright: noneisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Finding this balance becomes particularly difficult in the Java context, where software execution undergoes an initial warm-up phase due to just-in-time compilation (Barrett et al., 2017). During this phase, the Java Virtual Machine (JVM) performs a wide range of optimizations, leading to fluctuations in performance behavior (as shown in Fig. 1) that may adversely affect result quality (Georges et al., 2007; Traini et al., 2022). To mitigate this issue, it is common practice to conduct a number of “warm-up iterations,” with the sole goal of warming up the JVM, before starting to collect performance measurements.
@@
  • We make publicly available the source code of our framework and the pretrained TSC models to aid future studies [^1] Our replication package, including the dataset, source code and pre-trained TSC models, is publicly available at https://doi.org/10.5281/zenodo.13749258..
@@ 2.1. Java Microbenchmarking
Microbenchmarking is a type of unit-level performance testing, commonly used in Java context (Leitner and Bezemer, 2017). A microbenchmark repeatedly executes a small portion of code, such as a Java method, while gathering measurements of its execution time. Due to just-in-time compilation, microbenchmarks are subject to performance variability in the first phase of their execution,222 The JVM allows disabling JIT compilation to reduce performance variability. However, doing so would give an unrealistic representation of software performance, as JIT compilation is typically enabled in production environments (Oaks, 2014). also known as warm-up. During this phase, the JVM detects frequently executed loops or methods, and it dynamically compiles them into optimized machine code. After the completion of the warm-up phase, the microbenchmark is said to be executing at a steady-state of performance (Barrett et al., 2017).
@@
State-of-practice (SOP). Practitioners usually predefine a fixed number of warm-up iterations based on their domain expertise. To configure warm-up iterations, they typically rely on Java Microbenchmark Harness (JMH) (OpenJDK, 2014), the de-facto standard for building, configuring, and running Java microbenchmarks. JMH allows to configure the number of warm-up iterations directly in the microbenchmark’s source code, using Java annotations, as shown in Listing 1. Nevertheless, prior work has shown that using a fixed number of iterations may often produce suboptimal estimates of the warm-up phase (Georges et al., 2007; Laaber et al., 2020; Traini et al., 2022).
State-of-the-art (SOTA). Researchers have developed techniques that can dynamically stop warm-up iterations at run-time. Georges et al. (Georges et al., 2007) introduced a first such approach, using a preset threshold on the coefficient of variation (Everitt and Skrondal, 2010) to determine the end of the warm-up phase. However, subsequent studies have identified Georges et al.’s heuristic as both inaccurate (Kalibera and Jones, 2013) and unrealistic (Laaber et al., 2020; Traini et al., 2022) in practical scenarios. Another technique was recently proposed by Laaber et al. (Laaber et al., 2020). Similarly to Georges et al.’s heuristic, it leverages statistical metrics to estimate the end of the warm-up phase. At each microbenchmark iteration, this technique checks whether adding more performance measurements are likely to change the distribution of performance measurements, and it uses this information to dynamically stop warm-up iterations at runtime. Laaber et al. provide three different variants of this technique, based on distinct statistical metrics, namely coefficient of variation (COV) (Everitt and Skrondal, 2010), [^3] Both the approaches of Georges et al. (Georges et al., 2007) and Laaber et al. (Laaber et al., 2020) use the CV, but they apply it in different ways. Georges et al. use a fixed threshold on CV, whereas Laaber et al. verify whether the addition of more measurements changes the CV within a specified threshold. relative confidence interval width (Davison and Hinkley, 1997; Kalibera and Jones, 2012) (RCIW), and Kullback-Leibler divergence (Kullback and Leibler, 1951; He et al., 2019) (KLD). Prior work has shown that Laaber et al.’s technique is more accurate than the SOP in estimating the end of the warm-up phase (Traini et al., 2022).
@@ 3.4. Application
Our framework employs TSC models to dynamically predict when the microbenchmark reaches a steady-state of performance. As the microbenchmark execution progresses, the framework continuously evaluates the incoming performance measurements using the TSC model. Once the model detects the achievement of the steady-state, the framework promptly halts the microbenchmark execution and returns the current set of measurements for performance evaluation. Fig. 3 presents a snapshot of this procedure using an illustrative scenario. In this scenario, the microbenchmark has completed 3 warm-up iterations, with the current measurement window ranging from the 4th to the 7th iteration. Before proceeding, our framework queries the TSC model, forwarding the current set of measurements to determine if they correspond to a steady-state of execution. If the model predicts that the measurements are stable, the microbenchmark execution is halted, and these measurements are returned for performance evaluation. Conversely, if the model classifies them as unstable, then the framework makes the test run another iteration, and it shifts the measurement window by one iteration. This process repeats at each iteration until the model identifies a stable segment (or until a predefined maximum number of iterations is achieved).
@@ 5.1. RQ1: To what extent can TSC models accurately classify stable and unstable measurements?
Approach. We extract the number of warm-up iterations specified by the developers for each microbenchmark using a method similar to (Traini et al., 2022; Imran et al., 2024). We then compare the WEE of our framework with that of the SOP for each time series that reaches a steady-state. For this comparison, we employ the Wilcoxon signed-rank test (Wilcoxon, 1945) along with two measures of effect size: the Vargha-Delaney $\hat{A}_{12}$ (Vargha and Delaney, 2000) and the matched pairs rank biserial correlation $r$ (Kerby, 2014). In our context, the $\hat{A}_{12}$ measures the proportion of pairs where the WEE of our framework is lower than that of the developers. An $\hat{A}_{12}$ value $>0.5$ indicates that our framework provides a more accurate estimation of the warm-up phase than the SOP. The matched pairs rank biserial correlation $r$ represents the difference between the proportion of favorable and unfavorable evidence; in our case, favorable evidence indicates a lower WEE for the framework. Thus, an $r>0$ indicates that our framework performs better than the SOP.
@@ 6. Threats to validity
The proposed framework aims to dynamically infer the appropriate number of warm-up iterations at runtime. Although other relevant parameters, such as the number of forks and measurement iterations, can influence the quality of results and the testing time of microbenchmarks, these aspects are beyond the scope of our work. Nonetheless, to ensure fairness in our evaluation, we consistently use the same number of forks and measurement iterations when comparing our approach with baselines.
@@
Each TSC model is trained on an average of 417,520 segments from 468 microbenchmarks per fold iteration. Using training sets of different sizes and heterogeneity could affect the prediction accuracy of the models, and consequently, the outcome of the framework. We encourage future dedicated analyses to investigate how these factors might affect the framework effectiveness.
We utilize the dataset of JMH measurements from our previous work (Traini et al., 2022), which was collected using a rigorous procedure to minimize noise. We specifically choose these measurements to reduce confounding factors that could compromise the soundness of our study. However, using measurements from more variable environments (e.g., cloud) could lead to different study outcomes.
@@ 8. Conclusion
In this paper, we propose and study an AI-based framework to dynamically stop warm-up iterations in Java performance tests. The results show that, when integrated with certain TSC models, our framework delivers result quality comparable to the state-of-practice while drastically reducing the testing time. Furthermore, we find that the framework improves both the result quality and testing time in most of the investigated state-of-the-art techniques. This higher effectiveness can be attributed to the ability of TSC models to capture complex, non-linear patterns in measurements associated with steady-state execution, which simple statistical measures—such as those employed by state-of-the-art techniques–may overlook.